Find Equivalent processing time (min) at 100oC and 150oC, respectively, which will determine the lethality equivalent to that obtained at 121oC. F121 = 45min and Z= 10.5oC.
1
Expert's answer
2020-07-22T06:03:23-0400

For equivalent processing time we know that,

"F=F_{ref}\times 10^{\frac{T_{ref}-T}{z}}"


here it is given that, "F_{ref}=F_{121}=40" min, so for temp

(a) T=100oC,

"F=F_{ref}\times 10^{\frac{T_{ref}-T}{z}}"

"F=40\times 10^{\frac{121-100}{10.5}}"

F= 4000 min

(b) For temp T= 150 oC


"F=F_{ref}\times 10^{\frac{T_{ref}-T}{z}}"


"F=40\times 10^{\frac{121-150}{10.5}}"

F= 0.0767 min